### v1

# appAppendix References (eccv) Package eccv Warning: Package ‘hyperref’ is loaded with option ‘pagebackref’, which is *not* recommended for camera-ready version

Abstract
The diffusion model has long been plagued by scalability and quadratic complexity issues, especially within transformer-based structures. In this study, we aim to leverage the long sequence modeling capability of a State-Space Model called Mamba to extend its applicability to visual data generation. Firstly, we identify a critical oversight in most current Mamba-based vision methods, namely the lack of consideration for spatial continuity in the scan scheme of Mamba. Secondly, building upon this insight, we introduce a simple, plug-and-play, zero-parameter method named Zigzag Mamba, which outperforms Mamba-based baselines and demonstrates improved speed and memory utilization compared to transformer-based baselines. Lastly, we integrate Zigzag Mamba with the Stochastic Interpolant framework to investigate the scalability of the model on large-resolution visual datasets, such as FacesHQ $1024\times 1024$ and UCF101, MultiModal-CelebA-HQ, and MS COCO $256\times 256$. Code will be released at https://taohu.me/zigma/
Keywords:
Diffusion Model State-Space Model Stochastic Interpolant
1 Introduction
Figure 1: Motivation. Our Zigzag Mamba method improves the network’s position-awareness by arranging and rearranging the scan path of Mamba in a heuristic manner.
Diffusion models have demonstrated significant advancements across various applications, including image processing [68], video analysis [39], point cloud processing [88], and human pose estimation [28]. Many of these models are built upon Latent Diffusion Models (LDM)[68], which are typically based on the UNet backbone. However, scalability remains a significant challenge in LDMs[44]. Recently, transformer-based structures have gained popularity due to their scalability [66, 8] and effectiveness in multi-modal training [9]. Notably, the transformer-based structure DiT [66] has even contributed to enhancing the high-fidelity video generation model SORA [65] by OpenAI. Despite efforts to alleviate the quadratic complexity of the attention mechanism through techniques such as windowing [60], sliding [11], sparsification [16, 49], and hashing [17, 75], it remains a bottleneck for diffusion models.
On the other hand, State-Space Models [31, 34, 30] have shown great potential for long sequence modeling, competing with transformer-based methods. Several methods [29, 31, 26, 70] have been proposed to enhance the robustness [93], scalability [29], and efficiency [31, 32] of State-Space Models. Among these, a method called Mamba [29] aims to alleviate these issues through work-efficient parallel scanning and other data-dependent innovations. However, the advantage of Mamba lies in 1D sequence modeling, and extending it to 2D images is a challenging question. Previous works [98, 59] have proposed flattening 2D tokens directly by computer hierarchy such as row-and-column-major order, but this approach neglects Spatial Continuity, as shown in Figure 1. Other works [56, 62] consider various directions in a single Mamba block, but this introduces additional parameters and GPU memory burden. In this paper, we aim to emphasize the importance of Spatial Continuity in Mamba and propose several intuitive and simple methods to enable the application of Mamba blocks to 2D images by incorporating continuity-based inductive biases in images. We also generalize these methods to 3D with spatial-temporal factorization on 3D sequence.
In the end, Stochastic Interpolant [3] provides a more generalized framework that can uniform various generative models including, Normalizing Flow [14], diffusion model [71, 38, 73], Flow matching [54, 58, 4], and Schrödinger Bridge [55]. Previously, some works [63] explore the Stochastic Interpolant on relatively small resolutions, e.g., $256\times 256$, $512\times 512$. In this work, we aim to explore it in further more complex scenarios e.g., $1024\times 1024$ resolution and even in videos.
In summary, our contributions are as follows: Firstly, we identify the critical issue of Spatial Continuity in generalizing the Mamba block from 1D sequence modeling to 2D image and 3D video modeling. Building on this insight, we propose a simple, plug-and-play, zero-parameter paradigm named Zigzag Mamba (ZigMa) that leverages spatial continuity to maximally incorporate the inductive bias from visual data. Secondly, we extend the methodology from 2D to 3D by factorizing the spatial and temporal sequences to optimize performance. Secondly, we provide comprehensive analysis surrounding the Mamba block within the regime of diffusion models. Lastly, we demonstrate that our designed Zigzag Mamba outperforms related Mamba-based baselines, representing the first exploration of Stochastic Interpolants on large-scale image data ($1024\times 1024$) and videos.
2 Related Works
Mamba. Several works [84, 83] have demonstrated that the State-Space Model possesses universal approximation ability under certain conditions. Mamba, as a new State-Space Model, has superior potential for modeling long sequences efficiently, which has been explored in various fields such as medical imaging [90, 62, 87, 69], image restoration [33, 97], graphs [10], NLP word byte [81], tabular data [2], point clouds [52], and image generation [24]. Among them, the most related to us are VisionMamba[98, 59], S4ND [64] and Mamba-ND [51]. VisionMamba[98, 59] uses a bidirectional SSM in discriminative tasks which incurs a high computational cost. Our method applies a simple alternative mamba diffusion in generative models. S4ND [64] introduces local convolution into Mamba’s reasoning process, moving beyond the use of only 1D data. Mamba-ND [51] takes multi-dimensionality into account in discriminative tasks, making use of various scans within a single block. In contrast, our focus is on distributing scan complexity across every layer of the network, thus maximizing the incorporation of inductive bias from visual data with zero parameter burden.
Backbones in Diffusion Models. Diffusion models primarily employ UNet-based [38, 68] and ViT-based [66, 8] backbones. While UNet is known for high memory demands [68], ViT benefits from scalability [21, 15] and multi-modal learning [9]. However, ViT’s quadratic complexity limits visual token processing, prompting studies towards mitigating this issue [85, 19, 11]. Our work, inspired by Mamba [29], explores an SSM-based model as a generic diffusion backbone, retaining ViT’s modality-agnostic and sequential modeling advantages. Concurrently, DiffSSM [91] concentrates on unconditional and class conditioning within the S4 model [31]. DIS [24] mainly explores the state-space model on a relatively small scale, which is not the exact focus of our work. Our work significantly differs from theirs as it primarily focuses on the backbone design using the Mamba block and extends it to text conditioning. Furthermore, we apply our method to more complex visual data.
SDE and ODE in Diffusion models. The realm of Score-based Generative Models encompasses significant contributions from foundational works such as Score Matching with Langevin Dynamics (SMLD) by Song et al. [72] and the advent of Denoising Diffusion Probabilistic Models(DDPM) [38, 71]. These methodologies operate within the framework of Stochastic Differential Equations (SDEs), a concept further refined in the research of Song et al. [73]. Recent research strides, as exemplified by Karras et al. [45] and Lee et al. [50], have showcased the efficacy of employing Ordinary Differential Equation (ODE) samplers for diffusion SDEs, offering significant reductions in sampling costs compared to traditional approaches that entail discretizing diffusion SDEs. Furthermore, within the domain of Flow Matching methodologies [54, 57, 3, 4], both SMLD and DDPMs emerge as specialized instances under distinct paths of the Probability Flow ODE framework [73]. These models typically utilize velocity field parameterizations employing the linear interpolant, a concept that finds broader applications in the Stochastic Interpolant framework [3], with subsequent generalizations extending to manifold settings [12]. Explored in various topics such as image generation [54, 40], editing [43], video generation [20], motion generation [42], point cloud generation [88], Reinforcement Learning [96] and text generation [41]. The SiT model [63] scrutinizes the interplay between interpolation methods in both sampling and training contexts, albeit in the context of smaller resolutions such as $512\times 512$. Our research endeavors to extend these insights to a larger scale, focusing on the generalization capabilities for 2D images of $1024\times 1024$ and 3D video data.
3 Method
In this section, we begin by providing background information on State-Space Models [31, 34, 30], with a particular focus on a special case known as Mamba [29]. We then highlight the critical issue of Spatial Continuity within the Mamba framework, and based on this insight, we propose the Zigzag Mamba. This enhancement aims to improve the efficiency of 2D data modeling by incorporating the continuity inductive bias inherent in 2D data. Furthermore, we design a basic cross-attention block upon Mamba block to achieve text-conditioning. Subsequently, we suggest extending this approach to 3D video data by factorizing the model into spatial and temporal dimensions, thereby facilitating the modeling process. Finally, we introduce the theoretical aspects of stochastic interpolation for training and sampling, which underpin our network architecture.
3.1 Background: State-Space Models
State Space Models (SSMs) [31, 34, 30] have been proven to handle long-range dependencies theoretically and empirically [32] with linear scaling w.r.t sequence length. In their general form, a linear state space model can be written as follows:
$x^{\prime}(t)$ $=\mathbf{A}(t)x(t)+\mathbf{B}(t)u(t)$
$y(t)$ $=\mathbf{C}(t)x(t)+\mathbf{D}(t)u(t),$
mapping a 1-D input sequence $u(t)\in\mathbb{R}$ to a 1-D output sequence $y(t)\in\mathbb{R}$ through an implicit N-D latent state sequence $x(t)\in\mathbb{R}^{n}$. Concretely, deep SSMs seek to use stacks of this simple model in a neural sequence modeling architecture, where the parameters $\mathbf{A},\textbf{B},\mathbf{C}$ and $\mathbf{D}$ for each layer can be learned via gradient descent.
Recently, Mamba [29] largely improved the flexibility of SSMs by relaxing the time-invariance constraint on SSM parameters, while maintaining computational efficiency. By employing a work-efficient parallel scan, Mamba mitigates the impact of the sequential nature of recurrence, whereas fusing GPU operations removes the requirement to materialize the expanded state. In this paper, we focus on exploring the scanning scheme of Mamba in diffusion models to maximize the use of inductive-bias from multi-dimensional visual data.
Figure 2: ZigMa. Our backbone is structured in L layers, mirroring the style of DiT [66]. We use the single-scan Mamba block as the primary reasoning module across different patches. To ensure the network is positionally aware, we’ve designed an arrange-rearrange scheme based on the single-scan Mamba. Different layers follow pairs of unique rearrange operation $\Omega$ and reverse rearrange $\bar{\Omega}$, optimizing the position-awareness of the method.
3.2 Diffusion Backbone: Zigzag Mamba
DiT-Style Network. We opt to use the framework of ViT by AdaLN [66] rather than the skip-layer focused U-ViT structure [8], as ViT has been validated as a scalable structure in literature [65, 9, 15]. Considering the aforementioned points, it informs our Mamba network design depicted in Figure 4. The core component of this design is the Zigzag Scanning, which will be explained in the following paragraph.
Zigzag Scanning in Mamba. Previous studies [82, 91] have used bidirectional scanning within the SSM framework. This approach has been expanded to include additional scanning directions [56, 59, 92] to account for the characteristics of 2D image data. These approaches unfold image patches along four directions, resulting in four distinct sequences. Each of these sequences is subsequently processed together through every SSM. However, since each direction may have different SSM parameters ($\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$, and $\mathbf{D}$), scaling up the number of directions could potentially lead to memory issues. In this work, we investigate the potential for amortizing the complexity of the Mamba into each layer of the network.
Our approach centers around the concept of token rearrangement before feeding them into the Forward Scan block. For a given input feature $\mathbf{z}_{i}$ from layer $i$, the output feature $\mathbf{z}_{i+1}$ of the Forward Scan block after the rearrangement can be expressed as:
${\mathbf{z}}_{\Omega_{i}}$ $=\texttt{arrange}({\mathbf{z}}_{i},\Omega_{i}),$ (1)
$\bar{{\mathbf{z}}}_{\Omega_{i}}$ $=\texttt{scan}({\mathbf{z}}_{\Omega}),$ (2)
${\mathbf{z}}_{i+1}$ $=\texttt{arrange}(\bar{{\mathbf{z}}}_{\Omega_{i}},\bar{\Omega_{i}}),$ (3)
$\Omega_{i}$ represents the 1D permutation of layer $i$, which rearranges the order of the patch tokens by $\Omega_{i}$, and $\Omega_{i}$ and $\overline{\Omega}_{i}$ represent the reverse operation. This ensures that both $\mathbf{z}_{i}$ and $\mathbf{z}_{i+1}$ maintain the sample order of the original image tokens.
(a) sweep-scan
(b) zigzag-scan
(c) zigzag-scan with 8 schemes
Figure 3: The 2D Image Scan. Our mamba scan design is based on the sweep-scan scheme shown in subfigure (a). From this, we developed a zigzag-scan scheme displayed in subfigure (b) to enhance the continuity of the patches, thereby maximizing the potential of the Mamba block. Since there are several possible arrangements for these continuous scans, we have listed the eight most common zigzag-scans in subfigure (c).
Now we explore the design of the $\Omega_{i}$ operation, considering additional inductive biases from 2D images. We propose one key properties: Spatial Continuity. Regarding Spatial Continuity, current innovations of Mamba in images [98, 59, 56] often squeeze 2D patch tokens directly following the computer hierarchy, such as row-and-column-major order. However, this approach may not be optimal for incorporating the inductive bias with neighboring tokens, as illustrated in Figure 3. To address this, we introduce a novel scanning scheme designed to maintain spatial continuity during the scan process. Additionally, we consider space-filling, which entails that for a patch of size $N\times N$, the length of the 1D continuous scanning scheme should be $N^{2}$. This helps to efficiently incorporate tokens to maximize the potential of long sequence modeling within the Mamba block.
[…]
Figure 4: The Detail of our Zigzag Mamba block. The detail of Mamba Scan is shown in Figure 2. The condition can include a timestep and a text prompt. These are fed into an MLP, which separately modulates the Mamba scan for long sequence modeling and cross attention for multi-modal reasoning.
Deploying text-condition on Zigzag Mamba. While Mamba offers the advantage of efficient long sequence modeling, it does so at the expense of the attention mechanism. As a result, there has been limited exploration into incorporating text-conditioning in Mamba-based diffusion models. To address this gap, we propose a straightforward cross-attention block with skip layers built upon the Mamba block, as illustrated in Figure 4. This design not only enables long sequence modeling but also facilitates multi-token conditioning, such as text-conditioning. Furthermore, it has the potential to provide interpretability [13, 76, 37], as cross-attention has been utilized in diffusion models.
Generalizing to 3D videos by factorizing spatial and temporal information. In previous sections, our focus has been on the spatial 2D Mamba, where we designed several spatially continuous, space-filling 2D scanning schemes. In this section, we aim to leverage this experience to aid in designing corresponding mechanisms for 3D video processing. We commence our design process by extrapolating from the conventional directional Mamba, as depicted in Figure 5. Given a video feature input ${\mathbf{z}}\in\mathbb{R}^{B\times T\times C\times W\times H}$, we propose three variants of the Video Mamba Block for facilitating 3D video generation.
(a) Sweep-scan: In this approach, we directly flatten the 3D feature $\mathbf{z}$ without considering spatial or temporal continuity. It’s worth noting that the flattening process follows the computer hierarchy order, meaning that no continuity is preserved in the flattened representation.
(b) 3D Zigzag: Compared with we formulate the 2D zigzag in previous subsections, we follow the similar design to generalize it to 3D Zigzag to keep the continuity in 2D and 3D simultaneously. Potentially, the scheme has much more complexity. We heuristically list 8 schemes as well. However, we empirically find that this scheme will lead to suboptimal optimization.
(c) Factorized 3D Zigzag = 2D Zigzag + 1D Sweep: To address the suboptimal optimization issue, we propose to factorize the spatial and temporal correlations as separate Mamba blocks. The order of their application can be adjusted as desired, for example, "sstt" or "ststst", where "s" represents the spatial-zigzag Mamba and "t" represents the temporal-zigzag Mamba.
Computation Analysis. For a visual sequence $\mathbf{T}\in\mathbb{R}^{1\times\text{M}\times\text{D}}$, the computation complexity of global self-attention and $k$-direction mamba and our zigzag mamba are as follows:
$\zeta(\text{self-attention})=4\text{M}\text{D}^{2}+2\text{M}^{2}\text{D},$ (4)
$\zeta(\text{k-mamba})=k\times[3\text{M}(2\text{D})\text{N}+\text{M}(2\text{D})\text{N}^{2}],$ (5)
$\zeta(\text{zigzag})=3\text{M}(2\text{D})\text{N}+\text{M}(2\text{D})\text{N}^{2},$ (6)
where self-attention exhibits quadratic complexity with respect to sequence length M, while Mamba exhibits linear complexity (N is a fixed parameter, set to 16 by default). Here, $k$ represents the number of scan directions in a single Mamba block. Therefore, $k$-mamba and zigzag share linear complexity with respect to self-attention. Moreover, our zigzag method can eliminate the $k$ series, further reducing the overall complexity.
Upon completing the design of the Zigzag Mamba network for improved visual inductive-bias integration, we proceed to combine it with a new diffusion framework, as illustrated below.
Figure 5: The 3D Video Scan. (a) We illustrate the bidirectional Mamba with the sweep scan, where the spatial and temporal information is treated as a set of tokens with a computer-hierarchy order. (b) For the 3D zigzag-scan, we aim to maximize the potential of Mamba by employing a spatial continuous scan scheme and adopting the optimal zigzag scan solution, as depicted in Figure 3. (c) We further separate the reasoning between spatial and temporal information, resulting in a factorized combination of 2D spatial scan ($\Omega$) plus a 1D temporal scan ($\Omega^{{}^{\prime}}$) scheme.
3.3 Diffusion Framework: Stochastic Interpolant
Sampling based on vector ${\mathbf{v}}$ and score ${\mathbf{s}}$.
Following [3, 77], the time-dependent probability distribution $p_{t}({\mathbf{x}})$ of ${\mathbf{x}}_{t}$ also coincides with the distribution of the reverse-time SDE [6]:
$d{\mathbf{X}}_{t}={\mathbf{v}}({\mathbf{X}}_{t},t)dt+\frac{1}{2}w_{t}{\mathbf{s}}({\mathbf{X}}_{t},t)dt+\sqrt{w_{t}}d\bar{\mathbf{W}}_{t},$ (7)
where $\bar{\mathbf{W}}_{t}$ is a reverse-time Wiener process, $w_{t}>0$ is an arbitrary time-dependent diffusion coefficient, ${\mathbf{s}}({\mathbf{x}},t)=\nabla\log p_{t}({\mathbf{x}})$ is the score, and ${\mathbf{v}}({\mathbf{x}},t)$ is given by the conditional expectation
${\mathbf{v}}({\mathbf{x}},t)$ $=\mathbb{E}[\dot{\mathbf{x}}_{t}|{\mathbf{x}}_{t}={\mathbf{x}}],$ (8)
$=\dot{\alpha}_{t}\mathbb{E}[{\mathbf{x}}_{*}|{\mathbf{x}}_{t}={\mathbf{x}}]+\dot{\sigma}_{t}\mathbb{E}[{\boldsymbol{\varepsilon}}|{\mathbf{x}}_{t}={\mathbf{x}}],$
where $\alpha_{t}$ is a decreasing function of $t$, and $\sigma_{t}$ is an increasing function of $t$. Here, $\dot{\alpha}_{t}$ and $\dot{\sigma}_{t}$ denote the time derivatives of $\alpha_{t}$ and $\sigma_{t}$, respectively.
As long as we can estimate the velocity ${\mathbf{v}}({\mathbf{x}},t)$ and/or score ${\mathbf{s}}({\mathbf{x}},t)$ fields, we can utilize it for the sampling process either by probability flow ODE [73] or the reverse-time SDE (7). Solving the reverse SDE (7) backwards in time from ${\mathbf{X}}_{T}={\boldsymbol{\varepsilon}}\sim\mathcal{N}(0,\mathbf{I})$ enables generating samples from the approximated data distribution $p_{0}(\mathbf{x})\sim p({\mathbf{x}})$. During sampling, we can perform direct sampling from either ODE or SDEs to balance between sampling speed and fidelity. If we choose to conduct ODE sampling, we can achieve this simply by setting the noise term ${\mathbf{s}}$ to zero.
In [3], it shows that one of the two quantities ${\mathbf{s}}_{\theta}({\mathbf{x}},t)$ and ${\mathbf{v}}_{\theta}({\mathbf{x}},t)$ needs to be estimated in practice. This follows directly from the constraint
${\mathbf{x}}$ $=\mathbb{E}[{\mathbf{x}}_{t}|{\mathbf{x}}_{t}={\mathbf{x}}],$ (9)
$=\alpha_{t}\mathbb{E}[{\mathbf{x}}_{*}|{\mathbf{x}}_{t}={\mathbf{x}}]+\sigma_{t}\mathbb{E}[{\boldsymbol{\varepsilon}}|{\mathbf{x}}_{t}={\mathbf{x}}],$
which can be used to re-express the score ${\mathbf{s}}({\mathbf{x}},t)$ in terms of the velocity ${\mathbf{v}}({\mathbf{x}},t)$ as
${\mathbf{s}}({\mathbf{x}},t)$ $=\sigma_{t}^{-1}\frac{\alpha_{t}{\mathbf{v}}({\mathbf{x}},t)-\dot{\alpha}_{t}{\mathbf{x}}}{\dot{\alpha}_{t}\sigma_{t}-\alpha_{t}\dot{\sigma}_{t}}.$ (10)
Thus, ${\mathbf{v}}({\mathbf{x}},t)$ and ${\mathbf{s}}({\mathbf{x}},t)$ can be mutually conversed. We illustrate how to compute them in the following.
Estimating the score ${\mathbf{s}}$ and the velocity ${\mathbf{v}}$. It has been shown in score-based diffusion models [73] that the score can be estimated parametrically as ${\mathbf{s}}_{\theta}({\mathbf{x}},t)$ using the loss
$\mathcal{L}_{\mathrm{s}}(\theta)=\int_{0}^{T}\mathbb{E}[\|\sigma_{t}{\mathbf{s}}_{\theta}({\mathbf{x}}_{t},t)+{\boldsymbol{\varepsilon}}\|^{2}]{\mathrm{d}}t.$ (11)
Similarly, the velocity ${\mathbf{v}}({\mathbf{x}},t)$ can be estimated parametrically as ${\mathbf{v}}_{\theta}({\mathbf{x}},t)$ via the loss
$\mathcal{L}_{\mathrm{v}}(\theta)$ $=\int_{0}^{T}\mathbb{E}[\|{\mathbf{v}}_{\theta}({\mathbf{x}}_{t},t)-\dot{\alpha}_{t}{\mathbf{x}}_{*}-\dot{\sigma}_{t}{\boldsymbol{\varepsilon}}\|^{2}]{\mathrm{d}}t,$ (12)
where $\theta$ represents the Zigzag Mamba network that we described in the previous section, we adopt the linear path for training, due to its simplicity and relatively straight trajectory:
$\alpha_{t}=1-t,\qquad$ $\sigma_{t}=t.$ (13)
We note that any time-dependent weight can be included under the integrals in both (11) and (12). These weight factors play a crucial role in score-based models when $T$ becomes large [48, 47]. Thus, they provide a general form that considers both the time-dependent weight and the stochasticity.
4 Experiment
In this section, we begin by detailing the experimental setup concerning image and video datasets, as well as our training details. Subsequently, we delve into several in-depth analyses aimed at elucidating the rationale behind our method design across various resolutions. Finally, we present our results obtained from higher-resolution and more complex datasets.
4.1 Dataset and Training Detail
Image Dataset. To explore the scalability in high resolution, we conduct experiments on the FacesHQ $1024\times 1024$. The general dataset that we use for training and ablations is FacesHQ, a compilation of CelebA-HQ [89] and FFHQ [46], as employed in previous work such as [23, 25].
For text-conditioned generation, we conduct the experiments on the MultiModal-CelebA $256^{2}$,$512^{2}$ [89] and MS COCO $256\times 256$ [53] datasets. Both datasets are composed of text-image pairs for training. Typically, there are 5 to 10 captions per image in COCO and MultiModal-CelebA. We convert discrete texts to a sequence of embeddings using a CLIP text encoder [67] following Stable Diffusion [68]. Then these embeddings are fed into the network as a sequence of tokens.
Video dataset. UCF101 dataset is consists of 13,320 video clips, which are classified into 101 categories. The total length of these video clips is over 27 hours. All these videos are collected from YouTube and have a fixed frame rate of 25 FPS with the resolution of $320\times 240$. We randomly sample continuous 16 frames and resize the frames to $256\times 256$.
Training Details. We uniformly use AdamW [61] optimizer with $41e-$ 4 learning rate. For extracting latent features, we employ off-the-shelf VAE encoders. To mitigate computational costs, we adopted a mixed-precision training approach. Additionally, we applied gradient clipping with a threshold of 2.0 and a weight decay of 0.01 to prevent NaN occurrences during Mamba training. Most of our experiments were conducted on 4 A100 GPUs, with scalability exploration extended to 16 and 32 A100 GPUs. For sampling, we adopt the ODE sampling for speed consideration. For further details, please refer to the Appendix 9.3.
Table 1: Ablation of Scanning Scheme Number. We evaluate various zigzag scanning schemes. Starting from a simple ‘‘Sweep’’ baseline, we consistently observe improvements as more schemes are implemented.
MultiModal-CelebA256256256256 MultiModal-CelebA512512512512
[…]
Sweep 158.1 75.9 0.169 162.3 103.2 0.203
Zigzag-1 65.7 47.8 0.051 121.0 78.0 0.113
Zigzag-2 54.7 45.5 0.041 96.0 59.5 0.079
Zigzag-8 45.5 26.4 0.011 34.9 29.5 0.023
4.2 Ablation Study
Scan Scheme Ablation. We provide several important findings based on our ablation studies on MultiModal-CelebA dataset in various resolutions in Table 1. Firstly, switching the scanning scheme from sweep to zigzag led to some gains. Secondly, as we increased the zigzag scheme from 1 to 8, we saw consistent gains. This indicates that alternating the scanning scheme in various blocks can be beneficial. Finally, the relative gain between Zigzag-1 and Zigzag-8 is more prominent at higher resolutions ($512\times 512$, or longer sequence token number) compared to lower resolutions ($256\times 256$, or shorter sequence token number), this shows the great potential and more efficient inductive-bias incorporation in longer sequence number.
Spatial Continuity is Critical. We first explore the importance of spatial continuity in Mamba design by grouping patches of size $N\times N$ into various sizes: $2\times 2$, $4\times 4$, $8\times 8$, and $16\times 16$, resulting in groups of patch sizes $N/2\times N/2$, $N/4\times N/4$, $N/8\times N/8$, and $N/16\times N/16$, respectively. Then, we apply our designed Zigzag-8 scheme at the group level instead of the patch level. Figure 6 illustrates that with increased spatial continuity, notably improved performance is achieved. Furthermore, we compare our approach with random shuffling of $N\times N$ patches, revealing notably inferior performance under random shuffling conditions. All of these results collectively indicate that spatial continuity is a critical requirement when applying Mamba in 2D sequences.
Figure 6: Spatial Continuity Analysis. As we incrementally enlarge the patch group size, the continuous segment of the patch also expands. This enhances spatial continuity, which we find improves FID on MultiModal-CelebA 256, 512 dataset.
Ablation study about the Network and FPS/GPU-Memory. In Figure 7 (a,b),we analyze the forward speed and GPU memory usage while varying the global patch dimensions from $32\times 32$ to $196\times 196$. For the speed analysis, we report Frame Per Second (FPS) instead of FLOPS, as FPS provides a more explicit and appropriate evaluation of speed  [^1] https://github.com/state-spaces/mamba/issues/110#issuecomment-1916464012. For simplicity, we uniformly apply the zigzag-1 Mamba scan scheme and use batch size=1 and patch size=1 on an A100 GPU with 80GB memory. It’s worth noting that all methods share nearly identical parameter numbers for fair comparison. We primarily compare our method with two popular transformer-based Diffusion backbones, U-ViT [8] and DiT [66]. It is evident that our method achieves the best FPS and GPU utilization when gradually increasing the patching number. U-ViT demonstrates the worst performance, even exceeds the memory bounds when the patch number is 196. Surprisingly, DiT’s GPU utilization is close to our method, which supports our backbone choice of DiT from a practical perspective.
Furthermore, in Figure 7 (c) we conduct an ablation study on GPU memory and FPS among different variants of our method. We find that our method incurs nearly zero FPS and GPU memory burden when gradually increasing the Mamba Scan Schemes. This analysis provides insights into the comparative performance and efficiency of different diffusion backbones, highlighting the advantages of our proposed method.
Order Receptive Field. We propose a new concept in Mamba-based structure for non-1D data. Given that various spatially-continuous zigzag paths may exist in non-1D data, we introduce the term Order Receptive Field which denotes the number of zigzag paths explicitly employed in the network design.
Ablation study about the Order Receptive Field and FPS/GPU-Memory. As depicted in Fig. 10, Zigzag Mamba consistently maintains its GPU memory consumption and FPS rate, even with a gradually increasing Order Receptive Field. In contrast, our primary baseline, Parallel Mamba, along with variants like Bidirectional Mamba and Vision Mamba [59, 98], experience a consistent decrease in FPS due to increased parameters. Notably, Zigzag Mamba, with an Order Receptive Field of 8, can perform faster without altering parameters.
(a) FPS v.s. Patch Number.
(b) GPU Memory v.s. Patch Number.
(c) GPU usage of variants of our method.
(d) FPS v.s. Patch Size.
Figure 7: (a, b).GPU Memory usage and FPS between our method and transformer-based methods(U-VIT [8] and DiT [66]). (c). GPU memory usage and FPS among various variants of our method. (d). The relationship between patch size and FID under various resolutions.
(a) Order Receptive Field v.s. FPS.
(b) Order Receptive Field v.s. GPU Memory.
Figure 8: The ablation study about Order Receptive Field, FPS, GPU Memory.
Patch size. We conducted an ablation study on patch sizes ranging from 1, 2, 4, to 8 in Figure 7 (d), aiming to explore their behaviors under the framework of Mamba. The results reveal that the FID deteriorates as the patch size increases, aligning with the common understanding observed in the field of transformers [22, 79]. This suggests that smaller patch sizes are crucial for optimal performance.
Table 2: Main Results on MS-COCO dataset. Our method consistently outperforms the baseline and can achieve even better results when the training scale is increased.
[…]
Sweep 195.1
Zigzag-1 73.1
Bidirection Mamba [98] 60.2
Zigzag-8 41.8
Zigzag-8 $\times$ 16GPU 33.8
Table 3: Main result on FacesHQ-1024 dataset with 4,094 tokens in latent space. Our method can outperform the baseline and can achieve even better results when the training scale is increased.
[…]
Bidirection Mamba-16GPU [98] 51.1 66.3
Zigzag-Mamba -16GPU 37.8 50.5
Zigzag-Mamba -32GPU 26.6 31.2
4.3 Main Result
Main Result on 1024$\times$1024 FacesHQ. To elaborate on the scalability of our method within the Mamba and Stochastic Interpolant framework, we provide comparisons on a high-resolution dataset (1024$\times$1024 FacesHQ) in Table 3. Our primary comparison is against Bidirectional Mamba, a commonly used solution for applying Mamba to 2D image data [59, 98]. With the aim of investigating Mamba’s scalability in large resolutions up to 1,024, we employ the diffusion model on the latent space of $128\times 128$ with a patch size of 2, resulting in 4,096 tokens. The network is trained on 16 A100 GPUs. Notably, our method demonstrates superior results compared to Bidirectional Mamba. Details regarding loss and FID curves can be found in Appendix 9.2. While constrained by GPU resource limitations, preventing longer training duration, we anticipate consistent outperformance of Bidirectional Mamba with extended training duration.
COCO dataset. To further compare the performance of our method, we also evaluate it on the more complex and common dataset MS COCO. We compare with the Bidirection Mamba as the baseline in Table 2. It should be noted that all methods share nearly identical parameter numbers for fair comparison. We trained all methods using 16 A100 GPUs. please check Appendix 9.3 for details. As depicted in Table 2, our Zigzag-8 method outperforms Bidirectional Mamba as well as Zigzag-1. This suggests that amortizing various scanning schemes can yield significant improvements, attributed to better incorporation of the inductive bias for 2D images in Mamba.
UCF101 dataset. In Table 4, we present our results on the UCF101 dataset, training all methods using 4 A100 GPUs, with further scalability exploration conducted using 16 A100 GPUs. We mainly compare our method consistantly with Bidirection Mamba [98]. For the choice of the 3D Zigzag Mamba, please refer to Appendix 9.3. For Factorized 3D Zigzag Mamba in video processing, we deploy the sst scheme for factorizing spatial and temporal modeling. This scheme prioritizes spatial information complexity over temporal information, hypothesizing that redundancy exists in the temporal domain. Our results consistently demonstrate the superior performance of our method across various scenarios, underscoring the intricacy and effectiveness of our approach.
Figure 9: Visualization of various resolutions on FacesHQ $1024\times 1024$ and MultiModal-CelebA $512\times 512$. Our generated samples present high fidelity across various resolutions.
Table 4: Video Scan Scheme on UCF101 dataset. Our method outperforms the baseline and can achieve even better results when the training scale is increased.
[…]
Bidirection Mamba [98] -4GPU 256.1 320.2
3D Zigzag Mamba -4GPU 238.1 282.3
Factorized 3D Zigzag Mamba -4GPU 216.1 210.2
Bidirection Mamba [98] -16GPU 146.2 201.1
Factorized 3D Zigzag Mamba -16GPU 121.2 140.1
Visualization. We demonstrate the image visualization of our best results on FacesHQ 1024 and MultiModal-CelebA 512 in Figure 9. For the visualization of videos, please refer to Appendix 9.1. It is evident that the visualization is visually pleasing across various resolutions, indicating the efficacy of our methods.
5 Conclusion
In this paper, we present the Zigzag Mamba Diffusion Model, developed within the Stochastic Interpolant framework. Our initial focus is on addressing the critical issue of spatial continuity. We then devise a Zigzag Mamba block to better utilize the inductive bias in 2D images. Further, we factorize the 3D Mamba into 2D and 1D Zigzag Mamba to facilitate optimization. We empirically design various ablation studies to examine different factors. This approach allows for a more in-depth exploration of the Stochastic Interpolant theory. We hope our endeavor can inspire further exploration in the Mamba network design. In the future, we plan to examine this network design for discriminative tasks.
6 Acknowledgement
We would like to thank Timy Phan, Yunlu Chen, Di Wu, Divin Yan for the extensive proofreading.